### v1

@@ 5.6 Qualitative Analysis
We conducted a qualitative analysis of our proposed method on three video sequences from LasHer, namely rightbike, darktreesboy, and ab__\__rightlowerredcup__\__quezhen, which contain four of the most common challenging scenarios: low illumination and illumination variation, which degrade the reliability of the visible modality; similar appearance, which is a common challenge for tracking tasks; and temperature crossover, which degrades the reliability of the thermal infrared modality. We also compared our method with four state-of-the-art RGB-T tracking methods: ViPT-deep, ProTrack, APFNet and HMFT.
@@
Figure 13 (a)-(d) shows the qualitative evaluation results. Our method’s tracking results are the most stable and accurate in low illumination and illumination variation scenarios, fully exploiting the potential of RGB-T tracking. In the similar appearance scenario, our method can still accurately locate and estimate the target boundary, while other methods fail due to interference from similar objects. In the temperature crossover scenario, which is a unique challenge for RGB-T tracking, our method is clearly the closest to the ground truth. The qualitative test results demonstrate the high robustness of M3PT in various common challenging scenarios.
5.7 Parameters Comparison
Table 5: The comparison of the foundation model, two similar methods and our method in terms of parameter size and performance. SOT stands for the foundation model, Params represents the model’s parameter size, and Tuned Params represents the number of parameters involved in fine-tuning.
Params Tuned Params LasHer
[…]
SOT 92.12M - 51.5 41.2
ViPT-shaw 92.73M 0.61M 59.0 47.6
[…]
ViPT-deep 92.96M 0.84M 64.8 52.2
[…]
M3PT 93M 0.88M 67.3 54.2
[…]
We verify the parameter-efficiency and effectiveness of our proposed method by comparing it with the RGB tracking foundation model and two best prompt methods ViPT-deep and ViPT-shaw. Table 5 shows that our method adds only 0.88M extra parameters on the foundation model, which are the only parameters that need to be fine-tuned during training. Compared with ViPT-deep, our method adds only 0.04M parameters, but the evaluation results on LasHer show that our method surpasses ViPT-deep by 2.5 and 2.0 percentage points on the PR and SR metrics respectively. The comparison result proves the parameter-efficiency and effectiveness of our method.
5.8 Ablation Study
Table 6: The ablation experiment results. In the table, the four abbreviations represent the four prompting strategies. UEPS: Uni-modal Exploration Prompt Strategy, MFPS: Middle Fusion Prompt Strategy, FEPS: Fusion-modal Enhancing Prompt Strategy, MSPS: Modality-aware and Stage-aware Prompt Strategy.
UEPS MFPS FEPS MSPS RGBT234
[…]
63.5 48.3
$\surd$ 74.9 56.5
$\surd$ 79.2 58.5
$\surd$ 82.7 60.7
$\surd$ $\surd$ $\surd$ 84.5 62.2
$\surd$ $\surd$ $\surd$ $\surd$ 85.9 63.4
 [^1] The red, blue, and green fonts indicate the first, second, and third places of each evaluation metric, respectively.
Our proposed method consists of four prompt strategies: Uni-modal Exploration Prompt Strategy, Middle Fusion Prompt Strategy, Fusion-modal Enhancing Prompt Strategy and Modality-aware and Stage-aware Prompt Strategy. To verify the effectiveness of these strategies, we conduct extensive ablation studies and analyses on RGBT234. In the experiments, we build a baseline tracker which does not contain any prompt strategies and is composed entirely of upstream parameters, with MFP replaced by an element-wise addition operation of dual-modal features. Based on the baseline, we explore the effectiveness of the prompt strategies we propose.
Uni-modal Exploration Prompt Strategy: We investigate the importance of the Uni-modal Exploration Prompt Strategy we propose. We apply this strategy alone to the baseline tracker and retrain the corresponding prompt word generation branch. As shown in Table 6, compared to the original tracker, the addition of the Uni-modal Exploration Prompt Strategy improves the PR and SR of the tracker on RGBT234 by 11.4 and 8.2 percentage points, respectively, which proves that this strategy significantly enhances the adaptability of the first-stage backbone of foundation modal to the two uni-modal modeling tasks, and assists the backbone to fully explore the modality-shared features and modality-independent features.
Middle Fusion Prompt Strategy: To evaluate the impact of the Middle Fusion Prompt Strategy we propose, we replace the element-wise addition operation of the modal features in the original tracker with MFP, and provide fusion feature prompts for the second-stage backbone. According to the results in Table 10, the model with the Middle Fusion Prompt Strategy and retrained surpasses the original tracker by 19.2 and 12.4 percentage points on the PR and SR metrics on RGBT234, respectively. This indicates that this prompt strategy can achieve adaptive selection of modal discriminative features, and provide reliable fusion feature prompts for the second-stage backbone.
Fusion-modal Enhancing Prompt Strategy: We also validate the proposed Fusion-modal Enhancing Prompting Strategy. Table 6 shows that this strategy boosts the PR and SR metrics of the original tracker on RGBT234 by 19.2 and 12.4 percentage points, respectively. This confirms that our strategy better adapts the second-stage backbone to the fusion-modal modeling and feature enhancement.
Modality-aware and Stage-aware Prompt Strategy: To evaluate the effectiveness and necessity of the Modality-aware and Stage-aware Prompt strategy we propose, we build a tracker that only contains the other three prompt strategies and a tracker that contains all four prompt strategies, and retrain them separately. According to the results in Table 6, although the collaboration of the three prompt strategies for different purposes has already achieved excellent performance on RGBT234, the Modality-aware and Stage-aware Prompt strategy still improves the PR and SR metrics by 1.4 and 1.2 percentage points, respectively. By storing only a small amount of parameters for the low-level fixed patterns within the modality, the strategy enables the foundation model to identify the current modality and sub-task faster and better, and achieves more robust tracking.
@@ 6 Conclusion
In this work, we propose a novel middle fusion framework for RGB-T tracking, which achieves a balance between performance and efficiency. We also design a new prompt tracking method based on this framework, using four flexible visual prompt strategies. These strategies not only maximize the use of upstream knowledge, but also fully exploit the potential of prompt learning in RGB-T tracking tasks. We evaluate our method on four challenging RGB-T tracking benchmarks, and show that it outperforms the current best prompt tracking method, and leads the existing RGB-T tracking methods, while maintaining a high efficiency. Moreover, we offer new insights for the exploration of visual prompt learning.

### v2

@@ 5.12 Qualitative Analysis
From LasHer, We conduct a qualitative analysis of our proposed method on three video sequences, which contain four common challenging scenarios: low illumination and illumination variation, which degrade the reliability of visible modality; similar appearance, which is a common challenge for tracking tasks; and temperature crossover, which degrades the reliability of thermal infrared modality. We compare our method with four state-of-the-art RGB-T tracking methods: ViPT-deep [28], ProTrack [27], APFNet [15] and HMFT [17].
@@
According to the qualitative comparison results in Figure 14 (a)-(d), our method’s tracking results are the most stable and accurate in low illumination and illumination variation scenarios, fully exploiting the potential of RGB-T prompt tracking. In the similar appearance scenario, our method can still accurately locate and estimate the target boundary, while other methods fail due to interference from similar objects. In the temperature crossover scenario, which is a unique challenge for RGB-T tracking, our method is clearly the closest to the ground truth. The qualitative test results demonstrate the high robustness of M3PT in various common challenging scenarios.
Figure 15: A failure case in an out-of-view scenario.
Furthermore, corresponding to our attribute-wise evaluation results, we present a tracking failure case, containing an out-of-view scenario, to better analyse the reason of failure. As depicted in Figure 15, when the target disappears from view, the tracker is constrained to re-localize, visually similar interfering objects across two modalities. This ultimately culminates in the accumulation of model errors and the subsequent failure of tracking.
@@ 6 Conclusion
In this work, we propose a novel prompt tracking method based on a middle fusion meta-framework and using four flexible visual prompt strategies. The four prompt strategies not only maximize the use of upstream knowledge, but also fully exploit the potential of prompt learning in RGB-T tracking tasks. Evaluation results on six recent challenging RGB-T tracking benchmarks demonstrate the superiority of our method over the current best prompt fine-tuning tracking methods, while also showcasing competitive performance compared to most full fine-tuning methods. Moreover, we offer new insights for the exploration of visual prompt learning. In the future, we plan to involve introducing motion cues or linguistic prompts to better address extreme scenarios. Furthermore, we intend to extend our parameter-efficient prompting approach to various other multi-modal video understanding tasks.